OpenStudy (anonymous):

Multiply these binomials. (3x - 4)(2x3 - 5x2)

OpenStudy (anonymous):

\[6x^4-23x^3+20x^2\]

OpenStudy (king):

6x^4-23x^3+20x^2

OpenStudy (anonymous):

6x4−23x3

OpenStudy (anonymous):

sorry 6x^4-23x^3+20x^2

OpenStudy (callisto):

(3x - 4)(2x3 - 5x2) = 3x (2x3 - 5x2) -4(2x3 - 5x2) = 6x^4 - 15x^3 - 8x^3 +20x^2 = 6x^4 - 23x^3 +20x^2 Understand?